Stellar Co. purchased a equipment on January 1, 2015, for $610,500. At that time, it was estimated that the equipment would have a 10-year life and no salvage value. On December 31, 2018, the firm’s accountant found that the entry for depreciation expense had been omitted in 2016. In addition, management has informed the accountant that the company plans to switch to straight-line depreciation, starting with the year 2018. At present, the company uses the sum-of-the-years’-digits method for depreciating equipment.

Prepare the general journal entries that should be made at December 31, 2018, to record these events. (Ignore tax effects.) (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)

Date

Account Titles and Explanation

Debit

Credit

Dec. 31, 2018

(To correct for the omission of depreciation expense in 2016.)

Dec. 31, 2018

(To record depreciation expense for 2018.)
Answer #1
Date Account Titles and Explanation Debit Credit
Dec. 31, 2018 Retained earnings 99900
Accumulated depreciation-Equipment 99900
(To correct for the omission of depreciation expense in 2016)
Dec. 31, 2018 Depreciation expense 44400
Accumulated depreciation-Equipment 44400
(To record depreciation expense for 2018)

Working:

Sum of the years' digits = 10 + 9 + 8 + 7 + 6 + 5 + 4 + 3 + 2 + 1 = 55

Depreciation for 2016 = $610500 x 9/55 = $99900

The change in the accounting method from sum-of-the-years'-digits method to straight-line method is a change in an accounting estimate due to a change in an accounting principle and is hence accounted for prospectively.

Depreciation for 2015 ($610500 x 10/55) 111000
Depreciation for 2016 99900
Depreciation for 2017 ($610500 x 8/55) 88800
Total depreciation up to 2018 299700
Cost of depreciable assets 610500
Less: Accumulated depreciation up to 2018 299700
Carrying value at January 1, 2018 310800
Less: Salvage value 0
Depreciable value at January 1, 2018 310800
Remaining useful life (10 - 3) 7
Depreciation for 2018 ($310800/7 years) 44400